OpenStudy (anonymous):

i just did this problem can you check to see if i did it right? ---- 15+3(2x-9)=60 ---12(2x-9)=60 --- 24x-216=60 --- 24x/24 276/24 --- than the answer would be x=12?

OpenStudy (ajprincess):

oops no. u did it wrong. 15+3(2x-9)=60 15+3*2x-3*9=60 15+6x-27=60 Can u do from this?

OpenStudy (anonymous):

or 60 - 27 my bad

OpenStudy (chihiroasleaf):

I think you did a mistake on the second step... consider that \[15+3\left( 2x-9 \right)=60\] \[3\left( 2x-9 \right)=60-15\] you can continue the next step... :) I think you can also solve it this way... :)

OpenStudy (ajprincess):

it will be 15+6x-27=60 15+6x-27+27=60+27

OpenStudy (anonymous):

nowww dont i have to simplify 15+6x before i divide?

OpenStudy (anonymous):

well before i add THEN divide

OpenStudy (ajprincess):

U hav 15+6x=87 Subtract 15 both sides. Then divide by 6 both sides.

OpenStudy (anonymous):

x=2?

OpenStudy (anonymous):

87-15=12 then 12/6 = 2

OpenStudy (ajprincess):

R u sure 87-15=12?

OpenStudy (anonymous):

72 woops

OpenStudy (ajprincess):

Nw can u find x?

OpenStudy (anonymous):

yess x=12

OpenStudy (ajprincess):

ya u r right.:)

OpenStudy (anonymous):

Okay thank you for this !!!

OpenStudy (anonymous):

appreciate it . [:

OpenStudy (ajprincess):

yw.:)